9. Q = L2 K. Calculate the elasticity of substitution

1
Expert's answer
2022-02-02T02:17:02-0500

Q = L2K

Elasticity of substitution ("\alpha") = "\frac{percentage change in K/L}{percentage change in MRTS}"


% change in K/L: = "{dIn(K\/L)}"

MRTS: ="\frac{MPL}{MPK}"


="\frac{LK}{L^{2}}"

= "\frac{K}{L}"


"\alpha" = "\frac{dIn(K/L)}{dIn(K/L)}"


"\alpha" = 1